OpenStudy (anonymous):

how do i solve this by substitution? 3x + 5y = 38 4x − 2y = 16

OpenStudy (anonymous):

im a little confused

OpenStudy (solomonzelman):

you can play around with the second equation. 4x - 2y = 16 subtract 4x from both sides - 2y = 16 - 4x divide both sides by -2 y = -8 + 2x or, you can re-write it as y = 2x - 8

OpenStudy (solomonzelman):

you can substitute `2x - 8` instead of y, into the first equation. equation, 3x+5y=38 substitution, 3x+5(2x-8)=38 and etc....

OpenStudy (anonymous):

so the answer is x=6,y=-4?

OpenStudy (anonymous):

x = −4, y = 6 x = 4, y = 6 x = 6, y = −4 x = 6, y = 4

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

please show your work. it really seems as though you are guessing.